Circular No. 8646

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005lw, 2005lx, 2005ly, 2005lz Additional supernovae have been discovered on unfiltered CCD images:

2005lw by R. R. Prasad and W. Li (LOSS/KAIT; cf. IAUC 8645); 2005lx by independently by Yoshimi Ichimura (Yoshimi-machi, Hiki-gun, Saitamaken, Japan, 0.28-m f/5.5 Schmidt-Cassegrain reflector; communicated via S. Nakano, Sumoto, Japan) and by T. Boles (Coddenham, England, 0.35m reflector; tabulated data below from Boles); and 2005ly and 2005lz by T. Puckett (0.60-m automated supernova patrol telescope; cf. IAUC 8645; with G. Sostero for 2005ly and with R. Gagliano for 2005lz).

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005lw Dec. 14.57 11^{h}08^{m}03.^{s}21 −12^{o}29′09′′.4 18.8 1′′.5 E, 6′′.5 S
2005lx Dec. 19.95 2 22 39.01 +28 15 07.1 18.1 24′′.8 W, 17′′.9 S
2005ly Dec. 21.12 1 23 28.84 +30 46 45.7 18.6 5′′.4 E, 18′′.2 S
2005lz Dec. 24.07 2 10 49.76 +34 58 58.2 17.8 13′′.8 E, 6′′.3 S

Additional KAIT magnitudes for 2005lw in IC 672: Jan. 20.48 UT, [19.5; Nov. 21.58, [18.5; Dec. 15.60, 18.4. SN 2005lw is a type-II supernova (cf. CBET 321). Additional unfiltered CCD magnitudes for 2005lx in IC 221: 2003 Oct. 30, [18.5 (K. Itagaki, Yamagata, Japan; via Nakano); 2005 Oct. 27 and Nov. 9, [19.5 (Boles); Dec. 19.667, 16.9 (Ichimura; measured by Nakano); 20.401, 17.0 (Ichimura); 20.417, 18.0 (R. and Y. Kushida, Kitakoma, Yamanashi, Japan; via Nakano); 20.596, 17.6 (K. Kadota, Ageo, Japan; via Nakano). Boles adds that nothing is visible at the location of 2005lx on Digitized Sky Survey images from 1989 (limiting red mag 21.0) and 1987 (limiting blue mag 20.5). Additional position end figures for 2005lx: Dec. 19.687, 38^{s}.8, 06′′ (offset from galaxy center: 27′′ west, 19′′.5 south; measured by Nakano from Ichimura’s jpeg image); Dec. 20.417, 38^{s}.95, 08′′.0 (R. Kushida, through clouds with a 0.40-m f/10 reflector; measured by Y. Kushida); Dec. 20.596, 39^{s}.00, 07′′.7 (Kadota, 0.25-m f/5 reflector). Additional unfiltered CCD magnitudes for 2005ly in UGC 934, Nov. 3, 10, 17, 24, and Dec. 10, [19.3–[19.9 (Puckett); 21.77, 17.8 (Sostero, 0.45-m reflector, Udine, Italy). SN 2005ly is a type-IIn supernova (cf. CBET 327). Additional unfiltered CCD magnitudes for 2005lz in UGC 1666: Sept. 11, [19.6 (Puckett); Dec. 25.11, 17.7 (J. Newton, Portal, AZ, 0.35-m reflector). Nothing was present at the position of 2005lz on images taken by Puckett on numerous dates from 1999 to 2005 (limiting mag ranging from 18.4 to 20.1). SN 2005lz is a type-Ia supernova (cf. CBET 337).

2005 December 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green